# Optical-vortex-pulse induced nonequilibrium spin textures in spin-orbit coupled electrons

### v1

@@ I Introduction
The combination of topologically structured light [59, 60, 61, 62] and spin-orbit-coupled electrons allows for developing ultrafast and tailor-made manipulation of spins [63, 64, 65, 66, 67, 68]. Recently, it has been reported that vector vortex beams, which are structured light with spatially variant polarization, can manipulate the electron spin texture in the GaAs/AlGaAs quantum well by imprinting their spatial helicity structure  [66]. Optical vortices are another type of topologically structured light characterized by OAM, yet their specific role in ultrafast spin manipulation remains unclear. Specifically, Laguerre-Gaussian beams having photon energy near the band gap of GaAs have been used to investigate photoinduced spin polarization in semiconductors [63]. However, light with OAM up to ±$\pm 5\hbar$ cannot detect the OAM dependence of spin polarization. This is because the OAM of light does not significantly change the optical selection rules for interband transitions when an azimuthal photon momentum is much smaller than the momentum scale of electrons. Recently, it has been demonstrated that the OAM of light can induce a nonlocal interaction between light and matter, resulting in the distinct OAM dependence of nonlocal photocurrent [65].
@@
In this paper, we examine the spin response of spin-orbit-coupled electron systems to optical vortices [Fig. 1(a)]. In particular, we focus on the influence of excitations within the conduction bands. This situation differs from previous experiments [63, 65] that stimulate optical interband transitions occurring beyond the band gap. We perform numerical calculations on the spin response using linear response theory, considering both the case involving only the Rashba SOI and the case of the equal strengths of the Rashba and Dresselhaus SOIs. Here we demonstrate that vortex beams can imprint characteristic spin textures reflecting the OAM of light [69, 40, 70, 71, 72], when either Rashba or Dresselhaus SOI is dominant. We also show that the resultant spin response yields a clear dependence on the OAM of light. The necessary strength of the AC electric field, $E_{0}$, for detection is estimated to be $E_{0}=O(1{\rm-}10~{}{\rm kV}/{\rm cm})$, based on the realistic SOI values in GaAs/AlGaAs heterostructures. In our calculations, we consider vortex beams in the THz frequency range [73, 74], whose energy is much smaller than the band gap in semiconductor-semiconductor heterostructures. However, our results are applicable to optical vortices within the infrared and visible frequency ranges when the materials have a large Fermi surface and strong SOIs, such as heavy metals.
@@ II.1 Model Hamiltonian and symmetries
In this paper, we examine the magnetic response of spin-orbit-coupled electron systems to irradiated vortex beams [Fig. 1(a)]. Let us specifically consider the interface of the semiconductor-semiconductor GaAs/AlGaAs heterostructure. At the interface, the inversion symmetry is broken, allowing the Rashba SOI to be odd in the wave vector ${\bm{k}}$. We also consider another type of SOI that stems from the bulk inversion asymmetry of the underlying crystal, which is known as Dresselhaus SOI.
Figure 1: (a) Schematic image of our setup. A pulsed vortex beam is irradiated to a 2DEG formed in the quantum well of the GaAs/AlGaAs heterostructure. (b,c) Fermi surfaces (thick curves) and spin polarization (arrows) of spin-orbit-coupled electrons on the Fermi surfaces: (b) $\beta_{\rm D}/\alpha_{\rm R}=0$ and (c) $\beta_{\rm D}/\alpha_{\rm R}=1$.
@@
${\bm{h}}({\bm{k}})=\alpha_{\rm R}\left(k_{y}\hat{\bm{x}}-k_{x}\hat{\bm{y}}\right)+\beta_{\rm D}\left(k_{x}\hat{\bm{x}}-k_{y}\hat{\bm{y}}\right),$ (3)
where the unit vectors $\hat{\bm{x}}$ and $\hat{\bm{y}}$ are defined as $\hat{\bm{x}}=(1,0,0)$ and $\hat{\bm{y}}=(0,1,0)$. Equation (2) is the Hamiltonian density for free electrons with the effective mass, $m_{\rm eff}$, and the chemical potential, $\mu$. The vector components, ${\bm{h}}({\bm{k}})$, represent the effective in-plane magnetic field induced by the SOIs, where the strengths of the Rashba and Dresselhaus SOIs are denoted by $\alpha_{\rm R}$ and $\beta_{\rm D}$, respectively. The Hamiltonian can model electrons residing in III–V semiconductor quantum wells [75]. Electrons are confined to the two-dimensional quantum well formed at the interface of a semiconductor-semiconductor heterostructure. Here we take into account only excitations occurring within the conduction band, disregarding excitations from the valence band beyond the band gap. The latter excitations are insignificant when the photon energy of optical vortices is considerably lower than the band gap.
@@ II.2 Spin response of 2DEGs
[…]
@@
[…]
@@ II.3 Vortex beams
Since the wavelength of THz vortex beams is significantly longer than the thickness of the two-dimensional quantum well at the interface, we can disregard the $z$-dependence of ${\bm{E}}$. The polarization vector is defined as $\hat{\bm{e}}_{\lambda}=\lambda(\hat{\bm{x}}+i\lambda\hat{\bm{y}})/\sqrt{2}$ with the SAM of light, $\lambda$, where $\lambda=\pm 1$ ($\lambda=0$) denote circularly polarized (linearly polarized) light. The spatial profile of $u_{p,m}$ at the focal plane ($z=0$) is then given by
$u_{p,m}({\bm{x}})=\left(\frac{\sqrt{2}\rho}{w_{0}}\right)^{|m|}e^{-\frac{\rho^{2}}{{w^{2}_{0}}}}L^{|m|}_{p}\left(\frac{2\rho^{2}}{w^{2}_{0}}\right)e^{im\theta},$ (16)
where $L^{m}_{p}$ is the associated Laguerre polynomial with $p,m\in\mathbb{Z}$ and $w_{0}$ is the radius of the Gaussian beam. We have introduced the cylindrical coordinate with $\rho=\sqrt{x^{2}+y^{2}}$ and $\theta=\tan^{-1}(y/x)$. The wavefront of the vortex beams, or the isophase plane, takes a spiral shape around the propagation ($z$) axis. The single-valuedness of the electromagnetic fields requires $|u_{p,m}({\bm{x}})|$ to be zero at the center of the beam, resulting in a doughnut-shaped intensity profile in the transverse plane. In momentum space, $|u_{p,m}({\bm{q}})|$, the intensity of vortex beams also shows a doughnut shape and peaks around $|{\bm{q}}|\sim 1/w_{0}$. The radius $|{\bm{q}}|$ increases with $|m|$, indicating that for vortex beams with $m\neq 0$, nonzero $|{\bm{q}}|$ contributes to the spin-current response function.
@@
×$\times\exp\left[-\left(\frac{t-t_{0}}{\sigma_{E}}\right)^{2}-i\Omega t\right]\hat{\bm{e}}_{s}\bigg{\}},$ (17)
@@
Figure 2: Linear response of the spin density, ${\bm{S}}({\bm{x}},t)/S_{0}$, induced by monochromatic vortex beams with $(\lambda,m)=(1,0)$ (a1), $(1,1)$ (a2), $(1,2)$ (a3), $(-1,0)$ (c1), $(-1,1)$ (c2), and $(-1,2)$ (c3), where $\lambda$ and $m$ are the SAM and OAM of light, respectively. The scaled quantity, $S({\bm{x}},t)/S_{0}$, with $S_{0}\equiv\hbar/2$ represents the spin density per $1~{}\mu{\rm m}^{2}$. We also plot the intensity of the electric field induced by the vortex beams with $(\lambda,m)=(1,0)$ (b1), $(1,1)$ (b2), $(1,2)$ (b3), $(-1,0)$ (d1), $(-1,1)$ (d2), and $(-1,2)$ (d3). Here we set $\alpha_{\rm D}k_{\rm F}=0.1~{}{\rm meV}$ and $\beta_{\rm D}=0$. The arrows in (a*) and (c*) correspond to the local spins $(S_{x},S_{y})/S_{0}$, and the color represents the amplitude of the local spin, $|{\bm{S}}(x,y)|/S_{0}$. The thick arrows in (b*) and (d*) show the direction of the electric field, ${\bm{E}}({\bm{x}})/|{\bm{E}}({\bm{x}})|$.
@@ III.1 Calculated systems and parameters
In this paper, we consider 2DEGs residing in semiconductor-semiconductor heterostructures. In such heterostructures, the strengths of the SOIs, $k_{\rm F}\alpha_{\rm R}$ and $k_{\rm F}\beta_{\rm D}$, are of the order of $1~{}{\rm meV}$ and less than the Fermi energy $\varepsilon_{\rm F}$. In the following numerical calculations, we set $k_{\rm F}\alpha=0.1~{}{\rm meV}$ [78, 79, 80], which corresponds to the parameters of the GaAs/AlGaAs heterostructures, $\varepsilon_{\rm F}=20~{}{\rm meV}$ and $m_{\rm eff}=0.067m_{\rm e}$ [81], where $m_{\rm e}$ is the free electron mass. The set of these parameters satisfies $k_{\rm F}\alpha_{\rm R}\ll\varepsilon_{\rm F}$. In Sec. III.2, we focus on the spin response of 2DEGs with purely Rashba-type SOI ($\beta_{\rm D}/\alpha_{\rm R}=0$). The spin response in the case of ${\rm SU}(2)$ symmetric SOI with $\alpha_{\rm R}=\beta_{\rm D}$ is discussed in Sec. III.3. We also fix $T/\varepsilon_{\rm F}=0.1$ and $\Gamma=0.1(\alpha_{\rm R}k_{\rm F})$ in the following calculations, where the latter satisfies the weak impurity concentration condition, $\Omega\tau=\hbar\Omega/2\Gamma\gg 1$. We would like to emphasize that the qualitative results are insensitive to the choice of these values.
@@ III.2 Rashba SOI: $\beta_{\rm D}/\alpha_{\rm R}=0$
where $\hat{\bm{n}}({\bm{x}},t)={\bm{E}}({\bm{x}},t)/|{\bm{E}}({\bm{x}},t)|$ is the unit vector pointing to the local electric field and $C$ is a closed path enclosing the singularity of the vortex beam at ${\bm{x}}={\bm{0}}$. As shown in Figs. 2(a2) and 2(a3), the local spins rotate in a clockwise direction once and twice, respectively, when moving counterclockwise around a circle of radius $w_{0}$. In a similar manner, the spins in Figs. 2(c2) and 2(c3) rotate counterclockwise once and twice, respectively. The winding number of the imprinted spin textures is given by substituting $\hat{\bm{n}}=\hat{\bm{S}}/|\hat{\bm{S}}|$ in Eq. (20), which coincides with that of the vortex beams, i.e., $w[{\bm{S}}]=w[{\bm{E}}]=-\lambda m$. This demonstrates that the OAM of light can be encoded into the spin textures of electrons through the SOI. This conclusion is also applicable to the case of purely Dresselhaus-type SOIs.
@@
So far, we have demonstrated that circularly polarized vortex beams can imprint the OAM of light onto the spin texture in 2DEGs with the Rashba SOI. The polarization of light plays an essential role in manipulating the spin texture. As shown in Eq. (18), the orientation of the local spins reflects the spatial profiles of the electric field generated by vortex beams. For linearly polarized light ($\lambda=0$), the electric field is aligned to the ±$\pm\hat{\bm{y}}$ directions, leading to the formation of spin domains alternatively oriented in the $+\hat{\bm{x}}$ and $-\hat{\bm{x}}$ directions.

### v2

@@ I Introduction
The combination of topologically structured light [59, 60, 61, 62] and spin-orbit-coupled electrons allows for developing ultrafast and tailor-made manipulation of spins [63, 64, 65, 66, 67, 68]. Recently, it has been reported that vector vortex beams, which are structured light with spatially variant polarization, can manipulate the electron spin texture in the GaAs/AlGaAs quantum well by imprinting their spatial helicity structure  [66]. Optical vortices are another type of topologically structured light characterized by OAM, yet their specific role in ultrafast spin manipulation remains unclear. Specifically, Laguerre-Gaussian beams having photon energy near the band gap of GaAs have been used to investigate photoinduced spin polarization in semiconductors [63]. However, light with OAM up to $\pm 5\hbar$ cannot detect the OAM dependence of spin polarization. This is because the OAM of light does not significantly change the optical selection rules for interband transitions when an azimuthal photon momentum is much smaller than the momentum scale of electrons. Recently, it has been demonstrated that the OAM of light can induce a nonlocal interaction between light and matter, resulting in the distinct OAM dependence of nonlocal photocurrent [65].
@@
In this paper, we examine the spin response of spin-orbit-coupled electron systems to optical vortices [Fig. 1(a)]. In particular, we focus on the influence of excitations within the conduction bands in the metallic states. This situation differs from previous experiments [63, 65] that stimulate optical interband transitions occurring beyond the band gap. We perform numerical calculations on the spin response using linear response theory, considering both the case involving only the Rashba SOI and the case of the equal strengths of the Rashba and Dresselhaus SOIs. We theoretically analyze pulse lasers rather than continuous waves because many experiments use pulsed light fields to observe laser-driven spin textures. Here we demonstrate that vortex beams can imprint characteristic spin textures reflecting the OAM of light [69, 40, 70, 71, 72], when either Rashba or Dresselhaus SOI is dominant. We also show that the resultant spin response yields a clear dependence on the OAM of light. The necessary strength of the AC electric field, $E_{0}$, for detection is estimated to be $E_{0}=O(1{\rm-}10~{\rm kV}/{\rm cm})$, based on the realistic SOI values in GaAs/AlGaAs heterostructures. In our calculations, we consider vortex beams in the THz frequency range [73, 74], whose energy is much smaller than the band gap in semiconductor-semiconductor heterostructures. However, our results are applicable to optical vortices within the infrared and visible frequency ranges when the materials have a large Fermi surface and strong SOIs, such as heavy metals.
We note that when calculating laser-pulse-driven real-space spin textures in a spin-orbit-coupled metal, summing the optical transitions of each atom is not efficient. This is because, in metallic systems, many electrons with different wave vectors near the Fermi surface are excited simultaneously, which is quite different from the case of semiconductors. Specifically, in the former, the wave-vector space picture is effective for computing the laser response, whereas in the latter, direct calculation in real space is possible. Therefore, we will examine the laser-driven spin response by combining the linear response theory in wave-vector space with numerical calculations. We demonstrate that these microscopic analyses can produce laser-driven spin textures in metallic systems at a quantitative level.
@@ II.1 Model Hamiltonian and symmetries
In this paper, we examine the magnetic response of spin-orbit-coupled electron systems to irradiated vortex beams [Fig. 1(a)]. As shown in Fig. 1(b), the intensity of the electric field associated with the vortex beam sharply peaks along its propagation axis within the Rayleigh range, $z_{\rm R}$, which is $O(100~\mu{\rm m})$ in the THz frequency range. For the further details, see Sec. II.3. Let us specifically consider the interface of the semiconductor-semiconductor GaAs/AlGaAs heterostructure, which is much thinner than the Rayleigh range. At the interface, the inversion symmetry is broken, allowing the Rashba SOI to be odd in the wave vector $\bm{k}$}. We also consider another type of SOI that stems from the bulk inversion asymmetry of the underlying crystal, which is known as Dresselhaus SOI.
Figure 1: (a) Schematic image of our setup. A pulsed vortex beam is irradiated to a 2DEG formed in the quantum well of the GaAs/AlGaAs heterostructure. (b) The spatial profile of the electric field generated by the vortex beam along the propagation ($z$) axis, $|{\bm{E}}(x,y=0,z)|$ (the upper panel) and $|{\bm{E}}(\rho_{\rm max},z)|$ (the lower pannel), where $(\rho,z)$ is the cylindrical coordinate and $\rho_{\rm max}$ is defined so that $E_{0}\equiv\max|{\bm{E}}(\rho,z=0)|$ at $\rho=\rho_{\rm max}$. The electric field sharply peaks within the Rayleigh range, $z_{\rm R}=O(100~\mu{\rm m})$. The inset is the electric field at the focal plane ($z=0$), where $x/w_{0},~y/w_{0}\in[-3,3]$. Here we consider the vortex beam carrying the orbital angular momentum, $m=2$. For the further details, see Sec. II.3. (c,d) Fermi surfaces (thick curves) and spin polarization (arrows) of spin-orbit-coupled electrons on the Fermi surfaces: (c) $\beta_{\rm D}/\alpha_{\rm R}=0$ and (d) $\beta_{\rm D}/\alpha_{\rm R}=1$.
@@
[…]
where the unit vectors $\hat{\bm{x}}$ and $\hat{\bm{y}}$ are defined as $\hat{\bm{x}}=(1,0,0)$ and $\hat{\bm{y}}=(0,1,0)$. Equation (2) is the Hamiltonian density for free electrons with the effective mass, $m_{\rm eff}$, and the chemical potential, $\mu$. The vector components, ${\bm{h}}({\bm{k}})$, represent the effective in-plane magnetic field induced by the SOIs, where the strengths of the Rashba and Dresselhaus SOIs are denoted by $\alpha_{\rm R}$ and $\beta_{\rm D}$, respectively. The Hamiltonian can model electrons residing in III-V semiconductor quantum wells [75]. Electrons are confined to the two-dimensional quantum well formed at the interface of a semiconductor-semiconductor heterostructure. Here we take into account only excitations occurring within the conduction band, disregarding excitations from the valence band beyond the band gap. The latter excitations are insignificant when the photon energy of optical vortices is considerably lower than the band gap.
@@ II.2 Spin response of 2DEGs
$S_{\mu}({\bm{x}},t)=\int dt^{\prime}\int d{\bm{x}}^{\prime}\Upsilon_{\mu\nu}(\bm{x}-{\bm{x}}^{\prime},t-t^{\prime})E_{\nu}(\bm{x}^{\prime},t^{\prime}).$ (11)
@@
$\chi_{\mu\nu}(q)$ $\equiv-\int d\tau\left\langle\hat{S}_{\mu}({\bm{q}},\tau)\hat{j}_{\nu}(-{\bm{q}},\tau)\right\rangle e^{i\omega_{m}\tau}$
@@ II.3 Vortex beams
Since the wavelength of THz vortex beams is significantly longer than the thickness of the two-dimensional quantum well at the interface, we can disregard the $z$-dependence of $\bm{E}$}. The polarization vector is defined as $\hat{\bm{e}}_{\lambda}=\lambda(\hat{\bm{x}}+i\lambda\hat{\bm{y}})/\sqrt{2}$ with the SAM of light, $\lambda$, where $\lambda=\pm 1$ ($\lambda=0$) denote circularly polarized (linearly polarized) light.
The spatial profile of $u_{p,m}$ for vortex beams is determined by solving the Helmholtz equation, $({\bm{\nabla}}^{2}+k^{2})u_{p,m}=0$, in the cylindrical coordinate, where $k\equiv\Omega/c$ with the speed of light, $c$. To solve the equation, we employ the paraxial approximation that assumes $u_{p,m}({\bm{x}})=\tilde{u}_{p,m}({\bm{x}})e^{ikz}$ to satisfy the conditions, $|\frac{\partial^{2}\tilde{u}}{\partial z^{2}}|\ll|\frac{\partial^{2}\tilde{u}}{\partial x^{2}}|$, $|\frac{\partial^{2}\tilde{u}}{\partial y^{2}}|$, and $|\frac{\partial^{2}\tilde{u}}{\partial z^{2}}|\ll 2k|\frac{\partial\tilde{u}}{\partial z}|$. The amplitude distribution changes slowly with distance $z$ when compared to variations of $\tilde{u}({\bm{x}})$ in the lateral direction. Then, we cacn remove the second-order derivative term and the Helmholtz equation reduces to
$\left({\bm{\nabla}}^{2}_{\perp}+2ik\frac{\partial}{\partial z}\right)\tilde{u}({\bm{x}})=0,$ (16)
where $\Delta_{\perp}^{2}\equiv\partial^{2}_{x}+\partial^{2}_{y}$. The Laguerre-Gaussian modes form a complete set of this equation, which are given by
$u_{p,m}({\bm{x}})=$ $\frac{1}{\sqrt{w(z)}}\left(\frac{\sqrt{2}\rho}{w(z)}\right)^{|m|}e^{-\frac{\rho^{2}}{w(z)^{2}}}L^{|m|}_{p}\left(\frac{2\rho^{2}}{w(z)^{2}}\right)e^{im\theta}$
$\times\exp\left(-\frac{ik\rho^{2}z}{z^{2}+z^{2}_{\rm R}}-i(2p+m+1)\chi(z)\right),$ (17)
where $L^{m}_{p}$ is the associated Laguerre polynomial. We have introduced the cylindrical coordinate with $\rho=\sqrt{x^{2}+y^{2}}$ and $\theta=\tan^{-1}(y/x)$. The beam waist $w_{0}$ determines the width, $w(z)=w_{0}\sqrt{1+z^{2}/z^{2}_{\rm R}}$, and the Rayleigh range of the beam, $z_{\rm R}=kw^{2}_{0}/2$ with $k=\Omega/c$, where $c$ is the speed of light. The Rayleigh range is the distance along the $z$ axis from the focal plane ($z=0$) at which the cross section of the beam becomes twice its minimum value, indicating the length scale of sharply peaked intensities along the $z$ axis [see Fig. 1(b)]. We have also introduced the Gouy phase $\chi(z)=\tan^{-1}(z/z_{\rm R})$. At the focal plane ($z=0$), the LG mode reduces to
$u_{p,m}(\rho,\theta)=$ $\frac{1}{\sqrt{w_{0}}}\left(\frac{\sqrt{2}\rho}{w_{0}}\right)^{|m|}e^{-\frac{\rho^{2}}{w^{2}_{0}}}L^{|m|}_{p}\left(\frac{2\rho^{2}}{w^{2}_{0}}\right)e^{im\theta}.$ (18)
The wavefront of the vortex beams, or the isophase plane, takes a spiral shape around the propagation ($z$) axis. In Figs. 2(b1-b3) and 2(d1-d3), we plot the intensity and the direction of the electric field induced by the vortex beams. For $m=0$ the beam reduces to the usual Gaussian beam whose intensity peaks at the center ($\rho=0$) [Fig. 2(b1)]. The single-valuedness of the electromagnetic fields requires $|u_{p,m}({\bm{x}})|$ to be zero at the center of the beam, resulting in a doughnut-shaped intensity profile in the transverse plane for $m\neq 0$. In momentum space, $|u_{p,m}({\bm{q}})|$, the intensity of vortex beams also shows a doughnut shape and peaks around $|{\bm{q}}|\sim 1/w_{0}$. The radius $|{\bm{q}}|$ increases with $|m|$, indicating that for vortex beams with $m\neq 0$, nonzero $|{\bm{q}}|$ contributes to the spin-current response function.
@@
$\times\exp\left[-\left(\frac{t-t_{0}}{\sigma_{E}}\right)^{2}-i\Omega t\right]\hat{\bm{e}}_{s}\bigg\},$ (19)
@@ III.1 Calculated systems and parameters
In this paper, we consider 2DEGs residing in semiconductor-semiconductor heterostructures. In such heterostructures, the strengths of the SOIs, $k_{\rm F}\alpha_{\rm R}$ and $k_{\rm F}\beta_{\rm D}$, are of the order of $1~{\rm meV}$ and less than the Fermi energy $\varepsilon_{\rm F}$. In the following numerical calculations, we set $k_{\rm F}\alpha=0.1~{\rm meV}$ [78, 79, 80], which corresponds to the parameters of the GaAs/AlGaAs heterostructures, $\varepsilon_{\rm F}=20~{\rm meV}$ and $m_{\rm eff}=0.067m_{\rm e}$ [81], where $m_{\rm e}$ is the free electron mass. The set of these parameters satisfies $k_{\rm F}\alpha_{\rm R}\ll\varepsilon_{\rm F}$. In Sec. III.2, we focus on the spin response of 2DEGs with purely Rashba-type SOI ($\beta_{\rm D}/\alpha_{\rm R}=0$). The spin response in the case of ${\rm SU}(2)$ symmetric SOI with $\alpha_{\rm R}=\beta_{\rm D}$ is discussed in Sec. III.3. We also fix $T/\varepsilon_{\rm F}=0.1$ and $\Gamma=0.1(\alpha_{\rm R}k_{\rm F})$ in the following calculations, where the latter satisfies the weak impurity concentration condition, $\Omega\tau=\hbar\Omega/2\Gamma\gg 1$. We note that the frequency spectrum of the pulsed field in Eq. (19) is sharply localized around $\Omega$ and the standard deviation $\sqrt{2}/\sigma_{E}$ is much smaller than $\Omega$ with the current parameter set. Therefore, the full frequency spectrum satisfies the weak impurity concentration. We would also like to emphasize that the qualitative results are insensitive to the choice of these values.
Figure 2: Linear response of the spin density, ${\bm{S}}({\bm{x}},t)/S_{0}$, induced by monochromatic vortex beams with $(\lambda,m)=(1,0)$ (a1), $(1,1)$ (a2), $(1,2)$ (a3), (1,0)(-1,0) (c1), (1,1)(-1,1) (c2), and (1,2)(-1,2) (c3), where $\lambda$ and $m$ are the SAM and OAM of light, respectively. The scaled quantity, $S({\bm{x}},t)/S_{0}$, with $S_{0}\equiv\hbar/2$ represents the spin density per $1~\mu{\rm m}^{2}$. We also plot the intensity of the electric field induced by the vortex beams with $(\lambda,m)=(1,0)$ (b1), $(1,1)$ (b2), $(1,2)$ (b3), (1,0)(-1,0) (d1), (1,1)(-1,1) (d2), and (1,2)(-1,2) (d3). Here we set $\alpha_{\rm D}k_{\rm F}=0.1~{\rm meV}$ and $\beta_{\rm D}=0$. The arrows in (a*) and (c*) correspond to the local spins $(S_{x},S_{y})/S_{0}$, and the color represents the amplitude of the local spin, $|{\bm{S}}(x,y)|/S_{0}$. The thick arrows in (b*) and (d*) show the direction of the electric field, ${\bm{E}}({\bm{x}})/|{\bm{E}}({\bm{x}})|$.
@@ III.2 Rashba SOI: $\beta_{\rm D}/\alpha_{\rm R}=0$
[…]
@@
So far, we have demonstrated that circularly polarized vortex beams can imprint the OAM of light onto the spin texture in 2DEGs with the Rashba SOI. The polarization of light plays an essential role in manipulating the spin texture. As shown in Eq. (20), the orientation of the local spins reflects the spatial profiles of the electric field generated by vortex beams. For linearly polarized light ($\lambda=0$), the electric field is aligned to the $\pm\hat{\bm{y}}$ directions, leading to the formation of spin domains alternatively oriented in the $\hat{\bm{x}}$ and $\hat{\bm{x}}$ directions.